Update at 09:59:11(BST): All appears to be working well since the reboot.
Update at 17:06:49(GMT): The server is back up and responding well once more. We'll continue to check this for any further issues.
Server16 has become unresponsive, so we're currently performing a reboot.
We expect the services to be back in a few minutes and we'll continue to monitor it once back.
23rd March 2012, at 5:02PM
28th March 2012, at 8:59AM
Update at 12:08:13(GMT): The caching issues should now be resolved and we are seeing mail flow as normal through the servers.
21st March 2012, at 9:44AM
21st March 2012, at 12:08PM
Update at 09:00:02(GMT): The issue has now been resolved with this server.The engineers have now resolved the issues that were causing the service failures. All services are back up and running again and under monitoring.
The server k9 appears be having trouble with the main Apache and MySQL services which will stop websites and email from working.
Our engineers have been notified and are investigating now. Please accept our apologies for any inconvenience caused during this time.
16th March 2012, at 8:12AM
16th March 2012, at 9:00AM
Update at 14:46:48(GMT): The issue has now been resolved with this server.Reboot and update complete.
The server is undergoing a reboot and emergency maintenance . This was for security reasons. We apologise for any inconvenience caused.
14th March 2012, at 1:38PM
14th March 2012, at 2:46PM
Update at 12:45:20(GMT): We are no longer experiencing any delays on the sending of SMTP mail.
7th March 2012, at 2:53PM
8th March 2012, at 12:45PM
Update at 10:19:43(GMT): The mail queue on adam has now been resolved and mail for users on this server will be processed normally.
6th March 2012, at 10:10AM
8th March 2012, at 10:19AM
Update at 19:45:35(GMT): The issue has now been resolved with this server.Update at 19:29:21(GMT): All services should now be restored.
Update at 19:28:27(GMT): Our engineer has now restored most services. Any remaining issues will be addressed momentarily.
Update at 18:47:31(GMT): Our engineer is now on site to fix the last few remaining servers affected by this fault.
Update at 17:46:30(GMT): The following servers are still unavailable due to this network issue:
vps20
vps21
prohost10
prohost11
prohost12
prohost13
businesshost1
xenserve19
xenserve20
We have an engineer en route to the datacentre to fix this remaining set of servers now and will update this post again once this had been done.
5th March 2012, at 10:49AM
5th March 2012, at 7:45PM
Update at 09:50:41(GMT): The issue with this server is now resolved
3rd March 2012, at 2:44PM
5th March 2012, at 9:50AM
Update at 14:04:44(GMT): The mail queue on max has now been cleared and we are seeing email delivered to mailboxes.
2nd March 2012, at 10:05PM
12th March 2012, at 2:04PM
Update at 13:34:51(GMT): Queues on the SMTP array have returned to nominal levels and there shouldn't be any delays remaining connected with the use of this service.
29th February 2012, at 9:12AM
1st March 2012, at 1:34PM